# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v1

@@ 1.3. Contribution Summary
Given that NSVAD research has become an explicit hotspot in AI, computing, and IoT communities, and shows great potential for applications in emerging scenarios such as smart cities and mobile internet, we aim to provide a systematic and inspiring guidance. This article is aimed at researchers and engineers who understand the main concepts and basic knowledge of AI but have no experience in NSVAD. We provide a comprehensive statement of unsupervised, weakly supervised, and fully unsupervised VAD routes, as well as various types of emerging tasks to satisfy readers with different backgrounds and needs. The contributions of this article can be summarized in the following four points:
@@
Focusing on AD in IoVT from the NSAI perspective, we comprehensively sort out the UVAD, WsVAD, and FuVAD routes and state their basic assumptions, learning paradigms, applicability scenarios of each scheme.
We open source available resources (e.g., benckmark datasets, code bases, literatures, workshops, and tutorials) and provide our studies on NSAVD in industry and smart cities.
We analyze the developmental vein between various research routes by empirically reviewing the recent advances and discussing the future vision of NSVAD in the context of trends and concerns in NSAI and IoVT.
@@ 1.4. Section Navigation
Based on the relationship between the content of the individual sections and the NSVAD architecture shown in Fig. 1, we organize the remainder of this article and provide an intuitive navigation map shown in Fig. 3. Specifically, in Section 2, we state the general basics of VAD, including the task definition, type of anomalies, and application areas. Sections 35 elaborate on the learning paradigms and typical models of UVAD, WsVAD and FuVAD, respectively. We provide detailed explanations of classic methods for readers to understand the core ideas and implementations better. To assist readers to conduct research instantly, we provide a comprehensive introduction to existing datasets and evaluation metrics involved in the current work in Section 6. Section 7 presents our latest exploration of NSVAD, including how to develop high-performance VAD models for dynamic environments in industrial IoT and how to address arbitrary data bias in smart city scenarios. Finally, Section 8 discusses the future vision of NSVAD, providing an in-depth analysis of its existing challenges, development trends, and possible opportunities. This tutorial paper is intended for non-specialists, so we prioritize conceptual clarification and research horizon construction in the main papers. Content that may overlap with existing work (e.g., itemized introductions to reviewed papers, detailed explanations of classical methods, and comprehensive presentations of research case) is relegated to the Supplement.
@@ 2. Generalized Foundations of NSVAD
As a cross-cutting research topic, NSVAD has received attention from researchers from the fields of deep learning, video surveillance, mobile internet, and edge computing. In the early stage, VAD follows the problem definition and basic setting of conventional AD tasks, i.e., anomalous samples are regarded as outliers with different distributions (Ramachandra et al., 2020). The corresponding unsupervised methods (Cao et al., 2024; Han et al., 2024b; Ye et al., 2024) attempt to learn prototypical representations of regular events and consider test videos outside the distribution as anomalies. Facing the high-dimensional, complex background and entangled representational properties of video, researchers at this time, mainly with AD backgrounds, focused on introducing efficient video representation learning (e.g., Auto-Encoders (AEs) (Park et al., 2020), Generative Adversarial Networks (GANs) (Sun et al., 2024), Transformers (Feng et al., 2021b), and diffusion model (Basak and Gautam, 2024)) from the CV community.
With the emergence of large-scale video datasets and continuously updated high-performance GPUs, deep learning-driven Video Understanding (VU) technicals have made significant progress, facilitating the flow of VAD research to the cross-cutting topics of VU and AD tasks. Eventually, a series of new VAD routes, such as WsVAD with Multiple Instance Learning and FuVAD based on iterative learning. These methods depart from the open-world assumption under the unsupervised AD setting, i.e., that anomalies are varied and unbounded in real-world situations. Specifically, WsVAD introduces anomaly instances in the training set to facilitate the model’s understanding of the essential differences between negative and positive samples. Although WsVAD methods need to collect and label anomalies, abnormal events of interest in a given scenario are usually enumerable. WsVAD can output more reliable results for specific types of abnormal events because anomalous patterns are seen during the learning phase. FuVAD completely escapes the limitation of UVAD and WsVAD on training data and tries to learn anomaly detectors directly from raw videos, saving the data preparation cost and avoiding the negative effect of mislabeling.
@@
In recent years, the rise of multimodal learning (Xu et al., 2023; Ektefaie et al., 2023) and large multimodal models, such as CLIP (Radford et al., 2021; Wu et al., 2024c) and DALL-E (Reddy et al., 2021) has brought a new windfall for VAD. Inspired by this, researchers have proposed new tasks such as open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022) and open-vocabulary VAD (Wu et al., 2024b), indicating the trend of integration between VAD and generative AI research. We understand the cognitive differences resulting from the research backgrounds of these fields and view the emergence of new routes and tasks as a positive signal to promote VAD to systematic NSVAD research. In addition, NSVAD systems deployed in real-world scenarios must face the challenge of domain bias due to multi-view, cross-scenario videos and consider the limited storage and communication resources of end devices. Recent NSVAD advances (Liu et al., 2024b; Wu et al., 2020) have begun considering both algorithm optimization and model deployment to balance performance and overhead. This section aims to provide a unified concept definition for various WSVAD routes.
@@ 2.1. Related Terms
We first introduce AD-related terminologies, including Anomaly Detection (AD), Novelty Detection (ND), Open Set Recognition (OSR), and Outlier Detection (OD). It is worth mentioning that these terms are confusing for most researchers from the computer vision and IoT communities, and AD practitioners often struggle with the lack of unified conceptual explanations (Yang et al., 2021). To this regard, we follow the general community consensus and combine our experience to sort out the above terms, as shown in Fig. 4.
@@
AD aims to detect samples that deviate from the learned normality predefined by the training data (Samariya and Thakkar, 2023). Such deviations are usually classified into two categories: 1) covariate shift at the raw data level, i.e., label-independent distributional differences between semantically identical samples, such as visual effects due to differences in image styles, video scenes, and acquisition equipment parameters, and 2) semantic shift at the label level, i.e., samples belonging to different categories. The former, also known as Sensory Anomaly Detection (SenAD), aims at detecting samples with label-independent bias, and tasks associated with it include domain adaptation and domain generalization. Obviously, UVAD aims to detect anomalous samples with different semantic labels from the training set, termed Semantic AD (SemAD),.
ND is often confused with AD tasks due to the fact that its setting and goals are essentially the same as those of SemAD (Roady et al., 2020). Specifically, ND, also known as novelty class detection, aims to detect test samples that do not belong to any of the known training classes, suggesting that it is primarily concerned with semantic offsetting. It is worth noting that although ND is subdivided into single-class ND and multi-class AD based on the number of categories in the training samples, they are both modeled as binary classification problems, i.e., they only need to identify the unknown categories without focusing on the so-called secondary labels (Yang et al., 2021). In our opinion, the difference between ND and SemAD is that the former has an optimistic attitude towards unknown category data and does not regard them as erroneous as SemAD does. Thus, UVAD can also be referred to as video novelty detection.
OSR aims to train a Multi-Calss Classifier (MCC) to simultaneously accomplish fine-grained categorization of the in-distribution data while detecting the unknown data that only appear in the test phase, i.e., it requires traditional MCC to exclude novel categories automatically (Geng et al., 2020). None of the current VAD routes consider the specific categories of normal events, focusing only on how to discriminate abnormal samples. Although the training set usually contains different normal events, they are all simply labeled as 0. Although OSR and VAD are currently distinct in terms of technical solutions and objectives, complex multi-agent systems (e.g., autonomous cars) often require the model not only to detect arbitrary anomalies that have not been seen before but also to accurately categorize regular events for decision-making. The integration of OSR with VAD has given rise to the Open Vocabulary VAD (OVVAD).
@@
OD aims to detect significantly different samples from other observations, i.e., outliers (Boukerche et al., 2020). Compared to AD, ND, and OSR, where out-of-distribution samples are only visible in the testing phase, OD directly accepts all observations containing both in- and out-of-distribution samples as inputs in the training phase, which is the same as FuVAD. Therefore, compared to UVAD following the SemAD/ND task setting, FuVAD does not restrict the training data to include only regular instances but uses unfiltered videos to learn anomaly classifiers directly. Consequently, FuVAD is superior in dealing with the massive real-time video streams generated by large-scale IoVT systems in real world.
@@ 2.2. Definition and Type of Anomaly
NSVAD aims to automatically detect unusual events that may result in personal injury or financial loss. How anomalies are defined and what types are included are critical to conducting NSAVD research for real-world applications. As stated, UVAD and FuVAD follow the basic setups of SemAD and OD tasks, respectively. Thus, they define anomalies as a relative concept, i.e., samples distinguished from common data are all anomalies. In contrast, WsVAD departs from the open-set setting and focuses only on pre-defined specific kinds of anomalies, usually behaviors that humans intuitively perceive as abnormal. In Fig. 5, we show the composition of the training data for various routes.
Anomalies can be categorized into sensory anomalies focusing on domain offsets at the raw data level and semantic anomalies manifesting as semantic label distinctions. VAD aims at detecting anomalous events with positive semantic labels and is uninterested in label-independent offsets caused by the scene, acquisition device, and camera angle. Therefore, the anomalies defined in both UVAD and FuVAD fall into the category of semantic anomalies. Further, according to the information dimensional of the anomalies presented, we refer to (Ramachandra et al., 2020) to categorize them into appearance-only, motion-only, and joint appearance-motion anomalies, which correspond to deviation in spatial, temporal, and spatial-temporal interactions, respectively. Taking the CUHK Avenue dataset (Lu et al., 2013) as an example, the collector placed a red bag on the lawn, regarded as an appearance-only anomaly. Motion-only anomalies are usually similar to regular events in appearance but present incomprehensible motion, such as people wandering in public areas and cars traveling against the traffic flow. In the real world, anomalies usually present complexity and entanglement, not as single-dimensional information deviations but as misalignments in spatial-temporal interactions, termed appearance-motion united anomalies. It is difficult to effectively monitor such anomalies by relying on anomalous object detection or motion analysis alone. Therefore, the key to efficient UVAD is to make the model fully understand the prototypical features of regular events in appearance, motion, and spatial-temporal interactions (Liu et al., 2024b). Multi-proxy task-based UVAD models take this concern into account, aiming to utilize multiple optimization objectives to improve the model’s understanding of the difference between regular events and anomalous instances in different information dimensions.
In contrast, anomalies in the WsVAD dataset are more in line with human perceptions and real-world requirements, which are typically hazardous events that may pose a safety risk in a specific environment. For example, UCF-Crime (Sultani et al., 2018) collects 13 categories of common crimes against public safety, including abuse, arrest, arson, assault, road accidents, burglary, explosions, fights, robberies, shootings, theft, shoplifting, and vandalism. The XD-Vioence dataset (Wu et al., 2020) collects video and audio data of abuse, car accidents, explosions, fights, riots, and shootings. Although WsVAD violates open-world assumptions and fails to detect arbitrary anomalies outside of expectations, it has become a hot topic because of its reliable results on anomalies of human interest. Dataset producers usually predefine the anomalies according to the task requirements. Inspired by existing studies, WsVAD anomalies can be classified into short-term, long-term, and crowd anomalies. Compared with event labels, categorization according to the duration and manifestation of anomalies is more in line with WsVAD model development. Specifically, WsVAD mostly follows multiple instance learning, whereas earlier models assumed that the bags from anomaly videos contained only one anomaly segment, which is not valid for long-term anomalies. In addition, the magnitude of pattern differences in the exemplars formed from videos containing crowd anomalies is usually smaller than that of individual anomalies. The above two issues have been recognized as challenges that limit WsVAD (Liu et al., 2022b). In response, researchers have proposed Top- (Feng et al., 2021a), feature magnitude learning (Tian et al., 2021), and other improvement strategies. Therefore, how to make the model accurately correspond to more than one positive instance in long-term anomalies and perceive the weak differences between anomalies and normals in crowd anomalies is the key to developing high-performance WsVAD.
@@ 2.3. Definition of Various NSVAD Routes
UVAD refers to NSVAD schemes that use only easily collected routine events to train models to learn the spatial-temporal pattern boundaries of normal samples and to discriminate anomalies by measuring the distance between test samples and learned normality (Ramachandra et al., 2020). UVAD dominated early VAD research because it follows the open-world assumptions in the same vein as the AD community, circumventing predefinitions and collecting anomalous instances (Liu et al., 2021a). In Fig. 6, we show the general learning framework of UVAD and compare it with WsVAD and FuVAD. Specifically, UVAD assumes that models trained on regular events will only describe normal spatial-temporal patterns and will exhibit significant deviations when confronted with unseen anomalous examples, such as probability distributions (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), distances (Dalal et al., 2006; Saligrama and Chen, 2012; Dalal and Triggs, 2005), and proxy task errors (Liu et al., 2022c; Hasan et al., 2016; Gong et al., 2019; Liu et al., 2024b). Early approaches first used local binary operators (Hu et al., 2018; Nawarathna et al., 2014; Zhang et al., 2015), spatial-temporal points of interest (Dollár et al., 2005), etc. to characterize spatial-temporal features that are normal events, and then employed One-Class (OC) classifiers (e.g., OC support vector machines and OC neural networks) (Ionescu et al., 2019; Sun et al., 2019) to learn the pattern boundaries, and considered test samples whose features fell outside the boundaries as anomalous. Such methods rely on manual features and are prone to dimensionality disasters.
@@
In recent years, deep learning-driven UVAD integrates feature extraction and normality learning into a unified framework with two phases, training and testing, which corresponds to normality learning via the use of negative samples and anomaly detection via the detection of out-of-distribution samples (Liu et al., 2024c). In the normality learning phase, the network learnable parameters are optimized by minimizing a loss function over all negative samples. Whereas, in the testing phase, the degree of anomaly is measured by quantifying the distance between the test samples and the learned normality. Among them, reconstruction-based methods have dominated UVAD research in recent years (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Liu et al., 2022c, 2024b). On the one hand, most of the challenges faced by such methods, such as global motion modeling, temporal normality learning, and spatial detail inference, have been intensively studied in video self-supervised learning. As a result, many methods have driven the development of UVAD by drawing inspiration from existing methods such as video prediction (Oprea et al., 2020; Wang et al., 2022). On the other hand, since reconstruction/prediction methods aim to learn a generative model that can reason about regular events, their basic settings and optimization goals are clear and unambiguous, and thus easy to implement and follow. Essentially, UVAD is transductive learning, i.e., model training and testing are relatively independent. Numerous studies have shown that the performance of models in the normality learning phase on the agent task does not show a positive correlation with downstream anomaly detection. Due to the diversity of events, the spatial-temporal features of normal and abnormal samples overlap, and UVAD usually fails to actively recognize discriminative features and incorrectly learns some shared patterns when only regular events are available for model training (Liu et al., 2023c). In addition, Park et al. (Park et al., 2020) pointed out that overpowered deep neural networks may be able to effectively reason about unseen anomalous events during the testing phase due to overgeneralization performance, which may lead to underdetection. In response, the researchers proposed a memory network enhancement approach to weaken the model’s ability to generalize representations of anomalies by recording prototypical features.
@@
WsVAD uses weakly semantic video-level labels to supervise the output of strong semantic frame-level labels, i.e., frame-by-frame anomaly scores, by the sequencing model, thus enabling temporal localization of anomalous events. The first weakly-supervised approach is the multiple-example ranking framework proposed by Sultani et al. (Sultani et al., 2018) in 2018, which lays out the basic MIL architecture of the WsVAD route, and whose concurrently publicly available UCF-Crime dataset has become the most widely used weakly-supervised benchmark. They consider a video as a collection of multiple examples (video clips), where clips containing abnormal frames are positive examples, while clips with all normal frames are labeled as negative examples. Obviously, a normal video with label 0 produces all examples called negative bags, and the example-level labels are all . An abnormal video with the label constitutes a positive packet, which contains both positive and negative instances. Inspired by multiple-example learning, WsVAD aims to train a scoring model to output the anomaly scores of each example using video-level labels. The authors introduce a MIL ranking loss inspired by the hinge loss, which encourages the model to output high anomaly scores close to 1 for anomalous clips by maximizing the difference between the anomaly scores of the largest-scoring instances in the positive and negative bags while scoring regular clips as close to 0 as possible. In fact, WsVAD does not belong to any of the classes of out-of-distribution detection tasks introduced in Section 2.1, but is rather a type of multiple instance learning under weak semantic labeling supervision.
@@
Compared to UVAD, the weakly supervised approach introduces anomalous videos in the training set and provides video-level labels for all training samples. Although anomalous examples are diverse and unenumerable in the real world, noteworthy anomalous events in specific scenarios are usually limited and easy to collect, such as thefts, robberies, and traffic accidents, which can be obtained in large quantities from surveillance IoT systems and online video platforms. In addition, the labor cost of video-level tagging, which only requires marking whether a video contains an anomaly without worrying about the specific timing location (frame-level labeling) and detailed anomaly category, is usually affordable. For example, the UCF-Crime dataset is much larger than the one used by UVAD, but only 2100 discrete labels need to be provided. Due to the introduction of an additional human prior and the fact that the model has seen anomalous events during the training phase, WsVAD results are typically more reliable than UVAD, achieving excellent and consistent performance in detecting specific anomalous events. As a result, WsVAD has become a mainstream VAD scheme and is considered as the most promising research route for deployment in intelligent surveillance systems. The latest research attempts to mine anomaly-related clues from audio or subtitle text accompanying video frames, and proposes the multimodal WsVAD (Yang et al., 2024). This article introduces the recent progress of WsVAD in Section 4.
@@
FuVAD attempts to learn anomaly classifiers directly from large-scale raw videos without any editing and labeling (Pang et al., 2020; Zaheer et al., 2022). Specifically, FuVAD’s training data contains both positive and negative samples, and due to the low frequency of anomalous examples compared to regular events, the anomalous samples to be detected can be regarded as outliers with different patterns from the main data. In essence, the FuVAD model is transductive learning and does not follow the training-testing process of UVAD. Fully unsupervised methods have become a research hotspot in the internet era, where data preparation is costly, by virtue of the fact that they do not require any constraints and assumptions on the training data, and can be used to train the model by directly accessing huge amounts of videos from the real world (Liu et al., 2024c).

### v2

@@ 1.3. Contribution Summary
Given that NSVAD research has become an explicit hotspot in AI, computing, and IoT communities, and shows great potential for applications in emerging scenarios such as smart cities and mobile internet, we aim to provide systematic and inspiring guidance. This article is aimed at researchers and engineers who understand the main concepts and basic knowledge of AI but have no experience in NSVAD. We provide a comprehensive statement of unsupervised, weakly supervised, and fully unsupervised VAD routes, as well as various types of emerging tasks to satisfy readers with different backgrounds and needs. The contributions of this article can be summarized in the following four points:
@@
Focusing on AD in IoVT from the NSAI perspective, we comprehensively sort out the UVAD, WsVAD, and FuVAD routes and state their basic assumptions, learning paradigms, and applicability scenarios of each scheme.
We open source available resources (e.g., benchmark datasets, code bases, literature, workshops, and tutorials) and provide our studies on NSAVD in industry and smart cities.
We analyze the development sequence between various research routes by empirically reviewing the recent advances and discussing the future vision of NSVAD in the context of trends and concerns in NSAI and IoVT.
@@ 1.4. Section Navigation
Based on the relationship between the content of the individual sections and the NSVAD architecture shown in Fig. 1, we organize the remainder of this article and provide an intuitive navigation map shown in Fig. 3. Specifically, Section 2 states the general basics of VAD, including the task definition, type of anomalies, and application areas. Sections 35 elaborate on the learning paradigms and typical models of UVAD, WsVAD and FuVAD, respectively. We provide detailed explanations of classic methods for readers to understand the core ideas and implementations better. To assist readers to conduct research instantly, we provide a comprehensive introduction to existing datasets and evaluation metrics involved in the current work in Section 6. Section 7 presents our latest exploration of NSVAD, including how to develop high-performance VAD models for dynamic environments in industrial IoT and how to address arbitrary data bias in smart city scenarios. Finally, Section 8 discusses the future vision of NSVAD, providing an in-depth analysis of its existing challenges, development trends, and possible opportunities. This tutorial paper is intended for non-specialists, so we prioritize conceptual clarification and research horizon construction in the main papers. Content that may overlap with existing work (e.g., itemized introductions to reviewed papers, detailed explanations of classical methods, and comprehensive presentations of research cases) is relegated to the supplement.
@@ 2. Generalized Foundations of NSVAD
As a cross-cutting topic, NSVAD has attracted researchers from deep learning, video surveillance, mobile internet, and edge computing communities. Initially, VAD followed the conventional setting of AD problems, where anomalies were treated as outliers with different distributions (Ramachandra et al., 2020). Corresponding unsupervised methods (Cao et al., 2024; Han et al., 2024b; Ye et al., 2024) aimed to learn prototypical representations of regular events, considering test videos outside the distribution as anomalies. To address video’s high-dimensional and complex backgrounds, AD researchers introduced efficient video representation learning techniques like Auto-Encoders (AEs) (Park et al., 2020), Generative Adversarial Networks (GANs) (Sun et al., 2024; Huang et al., 2023; Wu et al., 2024a), Transformers (Feng et al., 2021b), and diffusion models (Basak and Gautam, 2024). With large-scale video datasets and high-performance GPUs, deep learning-driven Video Understanding (VU) techniques have advanced, shifting VAD research to cross-cutting topics of VU and AD. New VAD routes emerged, such as WsVAD with multiple instance learning (Sultani et al., 2018) and FuVAD with iterative learning (Pang et al., 2020). These methods challenge the open-world assumption under unsupervised AD, where real-world anomalies are varied and unbounded. WsVAD incorporates anomaly instances in training to differentiate between negative and positive samples. While WsVAD requires collecting and labeling anomalies, it outputs more reliable results for specific types of abnormal events. FuVAD avoids the data constraints of UVAD and WsVAD by learning anomaly detectors directly from raw videos, reducing data preparation costs and preventing mislabeling issues.
@@
In recent years, the rise of multimodal learning (Xu et al., 2023; Ektefaie et al., 2023) and large language models (Radford et al., 2021a; Reddy et al., 2021) have brought a new windfall for VAD. Researchers have proposed new tasks such as open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022), open-vocabulary VAD (Wu et al., 2024c), and Video Anomaly Retrieval (VAR) (Wu et al., 2024b), indicating the trend of integration between VAD and generative AI research. We understand the cognitive differences resulting from the research backgrounds of these fields and view the emergence of new routes and tasks as a positive signal to promote VAD to systematic NSVAD research. In addition, NSVAD systems deployed in real-world scenarios must face the challenge of domain bias due to multi-view, cross-scenario videos and consider the limited storage and communication resources of end devices. Recent NSVAD advances (Liu et al., 2024b; Wu et al., 2020) have begun considering both algorithm optimization and model deployment to balance performance and overhead.
@@ 2.1. Related Terms
We introduce key AD-related terms, including Anomaly Detection (AD), Novelty Detection (ND), Open Set Recognition (OSR), and Outlier Detection (OD). These terms often confuse researchers from the computer vision community (Wang et al., 2024a; Liu et al., 2024c), and AD practitioners struggle with inconsistent definitions (Yang et al., 2021). We follow community consensus and our experience to clarify these terms, as shown in Fig. 4.
@@
AD detects samples that deviate from normality as defined by training data (Samariya and Thakkar, 2023). Such deviations include: 1) covariate shift, i.e., label-independent distributional differences due to factors like image style or equipment, and 2) semantic shift, i.e., samples from different categories. The former is addressed by Sensory Anomaly Detection (SenAD), which includes tasks like domain adaptation. UVAD, by contrast, focuses on detecting anomalous samples with different semantic labels, termed Semantic AD (SemAD).
ND is often confused with AD since its goal is also to detect samples from unknown categories (Roady et al., 2020). ND is modeled as a binary classification problem, identifying unknown categories without concern for secondary labels (Yang et al., 2021). Unlike SemAD, ND views unknown data positively, making UVAD similar to video novelty detection.
OSR trains a Multi-Class Classifier (MCC) to categorize in-distribution data while detecting unknown data during testing (Geng et al., 2020). VAD generally does not categorize normal events but focuses on identifying anomalies. However, complex systems like autonomous cars require both anomaly detection and fine-grained categorization, leading to the integration of OSR and VAD in Open Vocabulary VAD (OVVAD).
@@
OD detects outliers, samples significantly different from others (Boukerche et al., 2020). Unlike AD, ND, and OSR, which detect out-of-distribution samples only during testing, OD accepts all data types during training, similar to FuVAD. FuVAD handles unfiltered videos to learn anomaly classifiers, making it superior for large-scale real-time video streams in IoVT systems.
@@ 2.2. Definition and Type of Anomaly
Defining anomalies and understanding their types is essential for real-world NSVAD applications. UVAD and FuVAD follow the setups of SemAD and OD tasks, where anomalies are relative, meaning anything differing from common data is considered an anomaly (Cheng et al., 2023b). In contrast, WsVAD focuses on specific pre-defined anomalies.
Specifically, anomalies are usually categorized as sensory (raw data deviations) or semantic (label differences). VAD targets semantic anomalies, ignoring irrelevant factors like the scene or camera angle changes. Anomalies in UVAD and FuVAD fall into appearance-only, motion-only, or appearance-motion categories, corresponding to deviations in spatial, temporal, or spatial-temporal interactions. For example, in the CUHK Avenue dataset (Lu et al., 2013), a red bag on a lawn is an appearance-only anomaly. More complex anomalies often involve misalignments in appearance-motion interactions, making them harder to detect with single-dimensional models. Therefore, effective UVAD models must understand regular event patterns in appearance, motion, and spatial-temporal contexts. Multi-proxy task-based models address this by improving the model’s ability to distinguish between normal and anomalous events across different dimensions. WsVAD focuses on real-world hazardous events, such as crimes in the UCF-Crime dataset (Sultani et al., 2018) and violent incidents in XD-Violence (Wu et al., 2020). Although WsVAD cannot detect arbitrary anomalies, its results are more reliable. WsVAD anomalies are often categorized as short-term, long-term, or crowd anomalies, aligning with real-world concerns.
@@ 2.3. Definition of Various NSVAD Routes
UVAD refers to NSVAD schemes that use only easily collected routine events to train models to learn the spatial-temporal pattern boundaries of normal samples (Ramachandra et al., 2020; Zhang et al., 2024a). UVAD dominated early VAD research because it follows the open-world assumptions in the same vein as the AD community, circumventing predefinitions and collecting anomalous instances (Liu et al., 2021a; Huang et al., 2024b). In Fig. 5, we show the general learning framework of UVAD and compare it with WsVAD and FuVAD. Specifically, UVAD assumes that models trained on regular events will only describe normal spatial-temporal patterns and will exhibit significant deviations when confronted with unseen anomalous examples, such as probability distributions (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), distances (Dalal et al., 2006; Saligrama and Chen, 2012; Dalal and Triggs, 2005), and proxy task errors (Liu et al., 2022c; Hasan et al., 2016; Gong et al., 2019; Liu et al., 2024b). Early approaches first used local binary operators (Hu et al., 2018; Nawarathna et al., 2014; Zhang et al., 2015), spatial-temporal points of interest (Dollár et al., 2005), etc. to characterize spatial-temporal features that are normal events, and then employed One-Class (OC) classifiers (e.g., OC support vector machines and OC neural networks) (Ionescu et al., 2019; Sun et al., 2019) to learn the pattern boundaries, and considered test samples whose features fell outside the boundaries as anomalous. Such methods rely on manual features and are prone to dimensionality disasters.
@@
In recent years, deep learning-driven UVAD integrates feature extraction and normality learning into a unified framework with two phases, training and testing, which corresponds to normality learning via the use of negative samples and anomaly detection via the detection of out-of-distribution samples (Liu et al., 2024d). In the normality learning phase, the network learnable parameters are optimized by minimizing a loss function overall negative samples. Whereas, in the testing phase, the degree of anomaly is measured by quantifying the distance between the test samples and the learned normality. Among them, reconstruction-based methods have dominated UVAD research in recent years (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Liu et al., 2022c, 2024b). On the one hand, most of the challenges faced by such methods, such as global motion modeling, temporal normality learning, and spatial detail inference, have been intensively studied in video self-supervised learning. As a result, many methods have driven the development of UVAD by drawing inspiration from existing methods, such as video prediction (Oprea et al., 2020; Wang et al., 2022). On the other hand, since reconstruction/prediction methods aim to learn a generative model that can reason about regular events, their basic settings and optimization goals are clear and unambiguous, and thus easy to implement and follow. Essentially, UVAD is transductive learning, i.e., model training and testing are relatively independent. Numerous studies have shown that the performance of models in the normality learning phase on the agent task does not show a positive correlation with downstream anomaly detection. Due to the diversity of events, the spatial-temporal features of normal and abnormal samples overlap, and UVAD usually fails to actively recognize discriminative features and incorrectly learns some shared patterns when only regular events are available for model training (Liu et al., 2023e). In addition, Park et al. (Park et al., 2020) pointed out that overpowered deep neural networks may be able to effectively reason about unseen anomalous events during the testing phase due to overgeneralization performance, which may lead to underdetection. In response, the researchers proposed a memory network enhancement approach to weaken the model’s ability to generalize representations of anomalies by recording prototypical features.
@@
WsVAD uses weakly semantic video-level labels to supervise the output of strong semantic frame-level labels, i.e., frame-by-frame anomaly scores, by the sequencing model, thus enabling temporal localization of anomalous events (Liu et al., 2022b; Karim et al., 2024). The first weakly-supervised approach is the multiple instance ranking framework proposed by Sultani et al. (Sultani et al., 2018) in 2018, which lays out the basic MIL architecture of the WsVAD route, and whose concurrently publicly available UCF-Crime dataset has become the most widely used weakly-supervised benchmark. They consider a video as a collection of multiple examples (video clips), where clips containing abnormal frames are positive examples, while clips with all normal frames are labeled as negative examples. Obviously, a normal video with label 0 produces all examples called negative bags, and the example-level labels are all . An abnormal video with the label constitutes a positive packet, which contains both positive and negative instances. Inspired by multiple instance learning, WsVAD aims to train a scoring model to output the anomaly scores of each example using video-level labels. The authors introduce a MIL ranking loss inspired by the hinge loss, which encourages the model to output high anomaly scores close to 1 for anomalous clips by maximizing the difference between the anomaly scores of the largest-scoring instances in the positive and negative bags while scoring regular clips as close to 0 as possible. In fact, WsVAD does not belong to any of the classes of out-of-distribution detection tasks introduced in Section 2.1, but is rather a type of multiple instance learning under weak semantic labeling supervision.
@@
Compared to UVAD, the weakly supervised approach introduces anomalous videos in the training set and provides video-level labels for all training samples. Although anomalous examples are diverse and unenumerable in the real world, noteworthy anomalous events in specific scenarios are usually limited and easy to collect, such as thefts, robberies, and traffic accidents, which can be obtained in large quantities from surveillance IoT systems and online video platforms. In addition, the labor cost of video-level tagging, which only requires marking whether a video contains an anomaly without worrying about the specific timing location (frame-level labeling) and detailed anomaly category, is usually affordable. For example, the UCF-Crime dataset is much larger than the one used by UVAD, but only 1,900 discrete labels need to be provided. Due to the introduction of an additional human prior and the fact that the model has seen anomalous events during the training phase, WsVAD results are typically more reliable than UVAD, achieving excellent and consistent performance in detecting specific anomalous events. As a result, WsVAD has become a mainstream VAD scheme and is considered as the most promising research route for deployment in intelligent surveillance systems. The latest research attempts to mine anomaly-related clues from audio or subtitle text accompanying video frames, and proposes the multimodal WsVAD (Yang et al., 2024b). This article introduces the recent progress of WsVAD in Section 4.
@@
FuVAD attempts to learn anomaly classifiers directly from large-scale raw videos without any editing and labeling (Pang et al., 2020; Zaheer et al., 2022). Specifically, FuVAD’s training data contains both positive and negative samples, and due to the low frequency of anomalous examples compared to regular events, the anomalous samples to be detected can be regarded as outliers with different patterns from the main data. In essence, the FuVAD model is transductive learning and does not follow the training-testing process of UVAD. Fully unsupervised methods have become a research hotspot in the internet era, where data preparation is costly, by virtue of the fact that they do not require any constraints and assumptions on the training data, and can be used to train the model by directly accessing huge amounts of videos from the real world (Liu et al., 2024d).